# Communication Cost Reduction for Subgraph Counting under Local Differential Privacy via Hash Functions

### v1

@@ VI-B2 Variance of the Modified Mechanism
When $j\neq j^{\prime}$, the two estimators are sampled from different sets of variables. Furthermore, the server sampling step is done independently for each group. The estimators are thus independent.
When $j=j^{\prime}$, $i\neq i^{\prime}$ and $k,k^{\prime}<j$, the server sampling step is done independently for $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i^{\prime})}_{j,h_{j}(k^{\prime})}$. Thus, by Theorem 5,
$\text{Cov}\left(\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i^{\prime})}_{j,h_{j}(k^{\prime})}\right)$$\leq$ $\text{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)$
$=$ $O\left(\frac{s^{2}}{n(e^{\epsilon_{1}}-1)^{2}}\right).$
Similarly, when $j=j^{\prime}$, $i\neq i^{\prime}$ and $k=k^{\prime}$, we have from Theorem 4 that:
$\text{Cov}\left(\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i^{\prime})}_{j,h_{j}(k)}\right)\leq\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)=O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\right).$
When $j=j^{\prime}$, $i=i^{\prime}$ and $k,k^{\prime}<j$, we will conduct the analysis by separating the case when $h_{j}(k)=h_{j}(k^{\prime})$ and the case when $h_{j}(k)\neq h_{j}(k^{\prime})$. Recall that $C_{k,k^{\prime}}^{(j)}$ is the event when $h_{j}(k)=h_{j}(k^{\prime})$. We have from Theorem 9 that:
$\mathbb{E}\left[\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}\mid C_{k,k^{\prime}}^{(j)}\right]$ $=$ $\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$
$=$ $O\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}\right).$
In the case where $k$ and $k^{\prime}$ are not colliding, we straightforwardly obtain from the proof of Theorem 5 that $\mathbb{P}\left(\overline{C^{(j)}_{k,k^{\prime}}}\right)\mathbb{E}\left[a^{(i)}_{j,h_{j}(k)}a^{(i)}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right]-\mathbb{E}\left[a^{(i)}_{j,h_{j}(k)}\right]\mathbb{E}\left[a^{(i)}_{j,h_{j}(k^{\prime})}\right]\leq 0$. Hence, we can say that:
$\text{Cov}\left(\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}\right)$
$\leq$ $\mathbb{P}\left(C^{(j)}_{k,k^{\prime}}\right)\mathbb{E}\left[\tilde{a}^{(i)}_{j,h_{j}(k)}\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right]$
$=$ $O\left(\frac{s^{2}}{\mu_{c}n(e^{\varepsilon_{1}}-1)^{2}}\right).$
Using the previous two theorems, we will quantify the variance of the triangle counting algorithm performed using GRR and central server sampling. We describe in the next theorem this variance prior to the addition of the Laplacian noise in the final step of the mechanism.
@@
Let $S_{2}$, $S_{3}$, $C_{4}$, and $W_{4}$ be the number of two-star, three-star, four-cycle, and the walk with length four in $G$. The variance of our estimator for triangle counting using the GRR and central server sampling technique, denoted by $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$, is $\mathcal{O}\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}\left(S_{2}+\mu_{c}C_{4}+\frac{\mu_{c}s}{n}W_{4}+\frac{s}{n}S_{3}\right)\right)$.
@@
The variance of $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$ comes from four parts:
@@
    $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)\right)$: We have from Theorem 9 that $\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$ is $O\left(\frac{s}{\mu_{c}(e^{e_{1}}-1)^{2}}\right)$. As there are $S_{2}$ terms in the summation, we have that the summation is no more than $O\left(\frac{s\cdot S_{2}}{\mu_{c}(e^{e_{1}}-1)^{2}}\right)$.
@@
    The covariance resulting from the repeated counting of $\{j,k\}$ in the summation: This is from the covariance of $a_{j,h_{j}(k)}^{(i)}$ and $a_{j,h_{j}(k)}^{(i^{\prime})}$ for $i,i^{\prime}>j>k$ such that $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{v_{i^{\prime}},v_{k}\}\in E$. By Theorem 10, the covariance of $a_{j,h_{j}(k)}^{(i)}$ and $a_{j,h_{j}(k)}^{(i^{\prime})}$ is $O\left(\frac{s}{(e^{e_{1}}-1)^{2}}\right)$, and the maximum frequency of such double counting is limited to no more than $C_{4}$.
@@
    The covariance between the terms $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i^{\prime})}_{j,h_{j}(k^{\prime})}$ for $i\neq i^{\prime}$ where both terms are included in the summation: This occurs when there are $i>j,k$ and $i^{\prime}>j,k^{\prime}$ such that the edge sets $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{v_{i^{\prime}},v_{k^{\prime}}\}$ are part of $E$. These instances occur no more than $W_{4}$ times. According to Theorem 5, each occurrence contributes to the covariance with a value of $\frac{s^{2}}{n}\cdot\frac{1}{(e^{e_{1}}-1)^{2}}$.
@@
    The covariance between the terms $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}$ where both terms are included in the summation: This occurs when there are $i>j,k$ such that the edge sets $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i},v_{k^{\prime}}\}$ are part of $E$. These instances occur no more than $S_{3}$ times. According to Theorem 5, each occurrence contributes to the covariance with a value of $\frac{s^{2}}{\mu_{c}n}\cdot\frac{1}{(e^{e_{1}}-1)^{2}}$.
@@
Let us compare the variance derived from the Generalized Random Response (GRR) as outlined in Theorem 7 with that obtained from GRR combined with central server sampling, as detailed in Theorem 11, particularly for graphs with numerous walks of length four. In such scenarios, the term $W_{4}$ becomes the dominant factor in the variances of both methods. We find that the constant associated with this dominant term is identical in both variances. This observation suggests that while achieving a similar level of variance, we can also reduce the download cost by a factor of $1/\mu_{c}$.
VI-C Smooth Sensitivity for Triangle Counting
As mentioned in the preceding subsection, the clipping method can be employed to diminish the magnitude of Laplacian noise introduced in the final stage of our mechanism. However, this method of clipping incurs a certain degree of bias in our estimator. To remove this bias, we suggest adopting the smooth sensitivity approach, which is introduced in Section II.
@@ VI-C Smooth Sensitivity for Triangle Counting
We now calculate the sensitivity at distance $k$ (defined in Definition 6) for the triangle counting. First, let us consider the local sensitivity (defined in Definition 5) for the adjacency vector $a^{\prime}\in\{0,1\}^{n}$ (denoted by $LS_{f}(a^{\prime})$). As previously discussed, the maximum change in the count $\sum\limits_{j,k:a^{\prime}_{j}=a^{\prime}_{k}=1}\tilde{a}^{(i)}_{j,h_{j}(k)}$ when one bit of $a^{\prime}$ is updated is no more than $UB_{\text{LS}}(a^{\prime})=\max\limits_{j}\left(\sum\limits_{k:a^{\prime}_{k}=1}\tilde{a}^{(i)}_{j,h_{j}(k)}+\sum\limits_{k:a^{\prime}_{k}=1}\tilde{a}^{(i)}_{k,h_{k}(j)}\right)$.
@@
When $a_{i}$ is the adjacent vector of user $i$ and $d_{i}$ is the number of ones in the vector, the number of ones in the vector $a$ with $|a-a_{i}|\leq k$ is not more than $UB_{\text{LS}}(a_{i})+k\cdot\omega_{i}$. Hence, the sensitivity at $k$ of the vector $a_{i}$, denoted by $A^{(k)}(a_{i})$ is no more than $O\left(UB_{\text{LS}}(a_{i})+\frac{k\cdot s}{e^{\varepsilon_{1}}+1}\right)$.
VI-D Experiments
Figure 1: Comparison on relative error between our algorithm and ARR for various download and upload costs on the Wikipedia Article Network graph with $\varepsilon=1$. In the legend, the abbreviation ’CSS’ denotes our Central Server Sampling method.
In this subsection, we aim to compare the algorithm we’ve outlined in earlier sections, which is fully accessible via our GitHub repository at https://github.com/Gericko/GroupRandomizedResponse, with the leading-edge algorithm named ARROne from [22]. It is worth noting that our implementation of ARROne differs slightly from the original version described in the cited paper, as we believe the original underestimates the sensitivity [43]. Specifically, in our version, the contribution of an edge $(v_{i},v_{j})$ to the count by user $v_{i}$ is defined as |$|\{k:a_{i,k}=1,\{v_{j},v_{k}\}\in M_{i},\text{ and }k<i\}|$ rather than |$|\{k:a_{i,k}=1,\{v_{j},v_{k}\}\in M_{i},\text{ and }j<k<i\}|$. Among the various algorithms discussed in [22], we focus on the one employing the 4-cycle trick, as it gives the highest performance.
Our experiments were performed on the Wikipedia Article Networks dataset [44, 45]. This dataset models Wikipedia users as nodes and their votes for admin promotions as edges. We selected this graph, which is based on real-world data, to assess the accuracy of our algorithm in practical scenarios. The Wikipedia dataset comprises 7,115 nodes and 100,762 edges. For experiments involving different graph sizes from the original, we created the extracted graph by randomly selecting a set of nodes of the desired size and then taking the subgraph induced by that set.
Our experiments focus on tracking the accuracy, quantified by relative error, as we modify various parameters. These parameters include the download cost, upload cost, size of the graph, and privacy budget.
The values presented for each parameter set represent the average derived from 10 separate simulation runs.
VI-D1 Error Analysis for Various Download Costs
The first experiment involves a comparison between the following three methods:
@@ VI-D1 Error Analysis for Various Download Costs
    The leading edge method detailed in [22]
@@
    Our GRR mechanism enhanced by the central server sampling and the clipping method [^1] The clipping method can also be applied in the counting step of the Generalized Random Response (GRR). We have established a theorem that specifies a clipping factor, ensuring that an edge is only clipped with a minimal probability. Nonetheless, we have opted not to delve into the details of the clipping factor in this paper, as our experiments indicate that the clipping method is outperformed by the application of smooth sensitivity. We have incorporated the results of the GRR mechanism, augmented with the clipping method, to verify and demonstrate its superiority over the ARR method.
  3. 3.
    Our GRR mechanism enhanced by the central server sampling and the smooth sensitivity

### v2

@@ 6.2.2 Variance of the Modified Mechanism
When $j\neq j^{\prime}$, the two estimators are based on disjoint sets of variables, and the server sampling step is conducted independently for each group. Therefore, the estimators are independent.
When $j=j^{\prime}$, $i\neq i^{\prime}$, and $k,k^{\prime}<j$, the sampling steps for $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i^{\prime})}_{j,h_{j}(k^{\prime})}$ are independent. Hence, by Theorem 5,
$\operatorname{Cov}\left(\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i^{\prime})}_{j,h_{j}(k^{\prime})}\right)\leq\operatorname{Cov}\left(\tilde{a}_{j,h_{j}(k)},\tilde{a}_{j,h_{j}(k^{\prime})}\right)=O\left(\frac{s^{2}}{n(e^{\varepsilon_{1}}-1)^{2}}\right).$
Similarly, when $j=j^{\prime}$, $i\neq i^{\prime}$, and $k=k^{\prime}$, Theorem 4 gives
$\operatorname{Cov}\left(\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i^{\prime})}_{j,h_{j}(k)}\right)\leq\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)=O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\right).$
Next, consider the case when $j=j^{\prime}$, $i=i^{\prime}$, and $k,k^{\prime}<j$. We distinguish two subcases depending on whether $h_{j}(k)=h_{j}(k^{\prime})$. Let $C^{(j)}_{k,k^{\prime}}$ denote the event that $h_{j}(k)=h_{j}(k^{\prime})$. By Theorem 9,
$\mathbb{E}\left[\tilde{a}^{(i)}_{j,h_{j}(k)}\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right]=\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)=O\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}\right).$
In the case $h_{j}(k)\neq h_{j}(k^{\prime})$, we obtain from the proof of Theorem 5 that
$\mathbb{P}\left(\overline{C^{(j)}_{k,k^{\prime}}}\right)\cdot\mathbb{E}\left[a^{(i)}_{j,h_{j}(k)}a^{(i)}_{j,h_{j}(k^{\prime})}\mid\overline{C^{(j)}_{k,k^{\prime}}}\right]-\mathbb{E}\left[a^{(i)}_{j,h_{j}(k)}\right]\mathbb{E}\left[a^{(i)}_{j,h_{j}(k^{\prime})}\right]\leq 0.$
Therefore,
$\operatorname{Cov}\left(\tilde{a}^{(i)}_{j,h_{j}(k)},\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}\right)\leq\mathbb{P}\left(C^{(j)}_{k,k^{\prime}}\right)\cdot\mathbb{E}\left[\tilde{a}^{(i)}_{j,h_{j}(k)}\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}\mid C^{(j)}_{k,k^{\prime}}\right]=O\left(\frac{s^{2}}{\mu_{c}n(e^{\varepsilon_{1}}-1)^{2}}\right).$
Using the previous two theorems, we will quantify the variance of the triangle counting algorithm performed using GroupRR and central server sampling. We describe in the next theorem this variance prior to the addition of the Laplacian noise in the final step of the mechanism.
@@
Let $S_{2}$, $S_{3}$, $C_{4}$, and $W_{4}$ be the number of two-star, three-star, four-cycle, and the walk with length four in $G$. The variance of our estimator for triangle counting using the GroupRR and central server sampling technique, denoted by $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$, is $\mathcal{O}\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}\left(S_{2}+\mu_{c}C_{4}+\frac{\mu_{c}s}{n}W_{4}+\frac{s}{n}S_{3}\right)\right)$.
@@
The variance of
$\sum_{i}\left(\sum_{\begin{subarray}{c}j,k:\\ \{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E\end{subarray}}\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$
can be decomposed into four components:
@@
    Variance of individual terms:
    $\sum_{i}\sum_{\begin{subarray}{c}j,k:\\ \{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E\end{subarray}}\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$
    By Theorem 9, we have
    $\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)=O\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}\right).$
    Since there are at most $S_{2}$ such terms in total, the overall contribution is $O\left(\frac{s\cdot S_{2}}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}\right)$.
@@
    Covariance from repeated index pairs $\{j,k\}$: This arises from the covariance between $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i^{\prime})}_{j,h_{j}(k)}$ for $i,i^{\prime}>j>k$ such that $\{v_{i},v_{j}\}$, $\{v_{i},v_{k}\}$, $\{v_{i^{\prime}},v_{j}\}$, and $\{v_{i^{\prime}},v_{k}\}$ are all in $E$. By Theorem 10, the covariance is bounded by $O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\right)$, and the number of such co-occurrences is at most $C_{4}$. Hence, this part contributes $O\left(\frac{s\cdot C_{4}}{(e^{\varepsilon_{1}}-1)^{2}}\right)$.
@@
    Covariance across different $i\neq i^{\prime}$: This refers to terms $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i^{\prime})}_{j,h_{j}(k^{\prime})}$ where $i\neq i^{\prime}$, and there exist $i>j,k$ and $i^{\prime}>j,k^{\prime}$ such that $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{v_{i^{\prime}},v_{k^{\prime}}\}\in E$. Each such occurrence contributes at most $O\left(\frac{s^{2}}{n(e^{\varepsilon_{1}}-1)^{2}}\right)$, and the total number of such cases is bounded by $W_{4}$, leading to a total contribution of $O\left(\frac{s^{2}\cdot W_{4}}{n(e^{\varepsilon_{1}}-1)^{2}}\right)$.
@@
    Covariance within the same $i$: This arises from pairs $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i)}_{j,h_{j}(k^{\prime})}$ for fixed $i$ and $j$, where $i>j$ and the edge set contains $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i},v_{k^{\prime}}\}\in E$. Each such pair contributes at most $O\left(\frac{s^{2}}{\mu_{c}n(e^{\varepsilon_{1}}-1)^{2}}\right)$, and there are at most $S_{3}$ such cases, giving a total contribution of $O\left(\frac{s^{2}\cdot S_{3}}{\mu_{c}n(e^{\varepsilon_{1}}-1)^{2}}\right)$.
@@
Let us compare the variance derived from the GroupRR as outlined in Theorem 7 with that obtained from GroupRR combined with central server sampling, as detailed in Theorem 11, particularly for graphs with numerous walks of length four. In such scenarios, the term $W_{4}$ becomes the dominant factor in the variances of both methods. We find that the constant associated with this dominant term is identical in both variances. This observation suggests that while achieving a similar level of variance, we can also reduce the download cost by a factor of $1/\mu_{c}$.
6.3 Smooth Sensitivity for Triangle Counting
@@
We calculate the sensitivity at distance $k$ (defined in Definition 6) for the triangle counting. First, let us consider the local sensitivity (defined in Definition 5) for the adjacency vector $a^{\prime}\in\{0,1\}^{n}$ (denoted by $LS_{f}(a^{\prime})$). As previously discussed, the maximum change in the count $\sum\limits_{j,k:a^{\prime}_{j}=a^{\prime}_{k}=1}\tilde{a}^{(i)}_{j,h_{j}(k)}$ when one bit of $a^{\prime}$ is updated is no more than
$UB_{\text{LS}}(a^{\prime})=\max\limits_{j}\left(\sum\limits_{k:a^{\prime}_{k}=1}\tilde{a}^{(i)}_{j,h_{j}(k)}+\sum\limits_{k:a^{\prime}_{k}=1}\tilde{a}^{(i)}_{k,h_{k}(j)}\right).$
@@
When $a_{i}$ is the adjacent vector of user $i$ and $d_{i}$ is the number of ones in the vector, the number of ones in the vector $a$ with $|a-a_{i}|\leq k$ is not more than $UB_{\text{LS}}(a_{i})+k\cdot\omega_{i}$. Hence, the sensitivity at $k$ of the vector $a_{i}$, denoted by $A^{(k)}(a_{i})$ is no more than $O\left(UB_{\text{LS}}(a_{i})+\frac{k\cdot s}{e^{\varepsilon_{1}}+1}\right)$. Combining with Theorem 11, we obtain that:
Theorem 12.
When using the smooth sensitivity mechanism, the variance of our estimator is $O(\frac{1}{\varepsilon_{2}^{2}}\sum\limits_{i=1}^{n}\left(UB_{\text{LS}}(a_{i})+\frac{s}{\varepsilon_{2}(e^{\varepsilon_{1}}+1)}\right)^{2}+\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}\left(S_{2}+\mu_{c}C_{4}+\frac{\mu_{c}s}{n}W_{4}+\frac{s}{n}S_{3}\right))$.
6.4 Clipping for Triangle Counting
Although the smooth sensitivity mechanism discussed in the previous subsection provides accurate results, calculating the sensitivity $UB_{\text{LS}}(a_{i})$ can be time-consuming in large networks. In such cases, we opt for the local Laplacian mechanism (Definition 3).
Let $f:\{0,1\}^{n}\rightarrow\mathbb{R}$ be defined as $f(a^{\prime})=\sum\limits_{j,k:a^{\prime}_{j}=a^{\prime}_{k}=1}\tilde{a}^{(i)}_{j,h_{j}(k)}$, representing the value we intend to publish for user $i$. Recall that the magnitude of the Laplacian noise added by the mechanism is determined by the global sensitivity $\Delta_{f}=\max\limits_{a^{\prime},a^{\prime\prime}:|a^{\prime}-a^{\prime\prime}|=1}\left(f(a^{\prime})-f(a^{\prime\prime})\right)$. Given that $\tilde{a}^{(i)}_{j,h_{j}(k)}\leq\omega_{i}$, we find that $\Delta_{f}\leq n\cdot\omega_{i}$.
We can use $n\cdot\omega_{i}$ as the magnitude of the Laplacian noise. However, $n\cdot\omega_{i}$ is to large that the noise can dominate the information we intend to publish. To have a better bound for the sensitivity, we employ the ideas of double clipping in Imola et al. (2022a). Although we use some ideas from the paper, as our mechanism is different from theirs, we attain a better sensitivity by a different mathematical analysis. Our double counting algorithm is as follows:
@@ 6.4 Clipping for Triangle Counting
    Degree Clipping: Recall that $\tilde{d}_{i}$ is the estimation for the degree of $v_{i}$ published in the first step of our mechanism. Let $\hat{d}_{i}=\tilde{d}_{i}+\varepsilon_{0}\ln\frac{2}{\beta}$ [^1] We notice that, in Imola et al. (2022a), edges are removed to match with $\hat{d}_{i}$ if $d_{i}>\hat{d}_{i}$. However, that process can be skipped, because we clip the sensitivity in the noisy triangle process anyway..
@@
    Noisy Triangle Clipping: Let Var represent the maximum variance of a single edge estimator, Cov the maximum covariance between two edge estimators, and define $b_{i}=\hat{d}_{i}+\sqrt{\frac{2}{\beta}(\hat{d}_{i}\cdot\text{Var}+\hat{d}_{i}^{2}\cdot\text{Cov})}$. Let $|\tilde{\Delta}_{i}^{(j^{\prime})}|=\sum\limits_{(j,k)\in W_{i}:j^{\prime}\in\{j,k\}}\tilde{a}^{(i)}_{j,h_{j}(k)}$, where $|\tilde{\Delta}_{i}^{(j^{\prime})}|$ denotes the contribution of node $j^{\prime}$ to $|\tilde{\Delta}_{i}|$. During the calculation of $|\tilde{\Delta}_{i}|$ in Line 4 of Algorithm 4, if there exists a $j^{\prime}$ such that $|\tilde{\Delta}_{i}^{(j^{\prime})}|>b_{i}$, we select a pair $(j,k)\in W_{i}$ such that $\tilde{a}^{(i)}_{j,h_{j}(k)}=\omega_{i}-\tilde{\sigma}_{i}$ and update the value to 0. This process is repeated until $|\tilde{\Delta}_{i}^{(j^{\prime})}|\leq b_{i}$.
The clipping process described above ensures that the contribution of any edge $(i,j^{\prime})$ to the publication of $|\tilde{\Delta}_{i}|$ does not exceed $b_{i}\ll n\cdot\omega_{i}$. As a result, the sensitivity of the publication of $|\tilde{\Delta}_{i}|$ is $b_{i}$, allowing us to set the Laplacian noise to this value.
While the clipping process allows us to have a smaller noise in the final step of our mechanism, the process can incur bias to our publication. We give a theoretical result for this clipping technique in the following theorem.
Theorem 13.
The clipping process incur no bias with probability at least $1-\beta$. The variance of our estimator by the clipping process is $O(\frac{\sum_{i}b_{i}^{2}}{\varepsilon_{2}^{2}}+\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}\left(S_{2}+\mu_{c}C_{4}+\frac{\mu_{c}s}{n}W_{4}+\frac{s}{n}S_{3}\right))$
We use the following 2 lemmas in our analysis.
Lemma 6.
The probability that we remove some edges in the degree clipping is smaller than $\beta/2$.
Proof.
We remove edges when $d_{i}>\hat{d}_{i}=\tilde{d}_{i}+\varepsilon_{0}\ln\frac{2}{\beta}$. This indicates that $\tilde{d}_{i}<d_{i}-\varepsilon_{0}\ln\frac{2}{\beta}$, meaning the magnitude of the Laplacian noise added during the degree-sharing step is greater than $\varepsilon_{0}\ln\frac{2}{\beta}$. This occurs with a probability smaller than $\beta/2$. ∎
While we consider the degree clipping in the previous lemma, we consider the noisy triangle clipping in the subsequent lemma.
Lemma 7.
The probability that we update some $\tilde{a}_{j,h_{j}(k)}^{(i)}$ in the noisy triangle clipping is not greater than $\beta/2$.
Proof.
By the degree clipping and the definition of $W_{i}$, the number of terms in the summation
$\sum_{(j,k)\in W_{i}:j^{\prime}\in\{j,k\}}\tilde{a}^{(i)}_{j,h_{j}(k)}$
is at most $\hat{d}_{i}$. From Theorem 8, we have that
$\mathbb{E}[\tilde{a}^{(i)}_{j,h_{j}(k)}]=a_{j,k}\in\{0,1\}.$
Therefore,
$\mathbb{E}\left[\sum_{(j,k)\in W_{i}:j^{\prime}\in\{j,k\}}\tilde{a}^{(i)}_{j,h_{j}(k)}\right]\leq\hat{d}_{i}.$
Since the summation contains at most $\hat{d}_{i}$ terms, its variance can be bounded by
$\text{Var}\left(\sum_{(j,k)\in W_{i}:j^{\prime}\in\{j,k\}}\tilde{a}^{(i)}_{j,h_{j}(k)}\right)\leq\hat{d}_{i}\cdot\text{Var}+\hat{d}_{i}^{2}\cdot\operatorname{Cov}.$
Using the bounds on the expectation and variance, we conclude that
$\Pr\left[|\tilde{\Delta}_{i}^{(j^{\prime})}|>b_{i}\right]\leq\frac{\beta}{2}.$
We are now ready to prove Theorem 13.
Proof of Theorem 13.
From Lemmas 6, 7, and union bound, we obtain that the clipping has no effect and incur no bias with probability at most $\beta/2$. The variance of the Laplacian noise added by the node $v_{i}$ is $b_{i}^{2}/\varepsilon_{2}^{2}$. Therefore, the final step of our mechanism increases the variance by $\sum_{i}b_{i}^{2}/\varepsilon_{2}^{2}$. ∎
In most practical cases, we have that $b_{i}=O\left(\frac{d_{i}\cdot s}{\sqrt{\beta}(e^{\varepsilon_{1}}-1)}\right)$. Based on that result, in all our experiments, the failure probability $\beta$ was set to $10^{-3}$.
One might wonder why we do not set the value of $b_{i}$ to $\max\limits_{j^{\prime}}|\tilde{\Delta}_{i}^{(j^{\prime})}|$, as it is smaller than the value used in the current clipping process. However, this is not feasible because $\max\limits_{j^{\prime}}|\tilde{\Delta}_{i}^{(j^{\prime})}|$ contains sensitive information, and we cannot determine the noise parameter of our mechanism based on such private data.
In contrast, the variance and covariance required for the computation of the clipping parameter $b_{i}$ can be computed using $n,s,\varepsilon$ and $d_{i}$. With the exception of $d_{i}$, all those are public information. Concerning $d_{i}$, it can be bounded with high probability by 0 and $\hat{d}_{i}$ to obtain a bound on $b_{i}$. Note that in the rare case that $d_{i}>\hat{d}_{i}$, the clipping mechanism still guaranties the privacy protection.
7 Experiments
In this section, we assess our method, GroupRR, in comparison to the leading communication-constrained graph publishing technique, ARR, as detailed in Imola et al. (2022a). As Imola et al. (2022a) has proven itself to be a more accurate algorithm than the rest of the state of the art, we restrict ourself to the comparison with this algorithm. We focus on the task of counting triangles in Subsection 7.1 and the task of counting 4-cycles in Subsection 7.2. All the code used for these experiments can be accessed at the following address https://github.com/Gericko/GroupRandomizedResponse.
7.1 Triangle Counting
(a) Wikipedia Article Network
(b) Facebook
Figure 3: Comparative analysis of relative error between our algorithm and ARR across different download and upload costs applied to the Wikipedia Article Network and the Facebook graphs, assuming a privacy budget of $\varepsilon=1$. The values of the parameter $s$ used for these experiments vary from 1 to 26 for the Wikipedia Article Network, and from 1 to 21 for Facebook, resulting in values of $\mu_{c}$ varying from 1 to $16,384$ and from 1 to $8,192$ respectively. The term ‘CSS’ in the legend stands for our central server sampling method
In this subsection, we compare the triangle counting algorithm described in earlier sections with the leading-edge algorithm named ARROne from Imola et al. (2022a) [^2] It is important to note, as mentioned in Appendix I of Imola et al. (2022a), that the original version of the algorithm underestimates the sensitivity. Consequently, we have adjusted the sensitivity calculations used in our evaluations. Specifically, in our modified version, the contribution of an edge $(v_{i},v_{j})$ to the count by user $v_{i}$ is defined as $|\{k:a_{i,k}=1,\{v_{j},v_{k}\}\in M_{i},\text{ and }k<i\}|$ instead of $|\{k:a_{i,k}=1,\{v_{j},v_{k}\}\in M_{i},\text{ and }j<k<i\}|$.. Among the various algorithms discussed in the paper, our analysis focuses on the one that utilizes the 4-cycle trick, which has been shown to yield the highest performance.
Our experiments primarily utilized the Wikipedia Article Networks dataset (Leskovec et al., 2010b; a), which contains 7,115 nodes and 100,762 edges, and the Facebook dataset (Leskovec & Mcauley, 2012), which contains 4,039 nodes and 88,234 edges. We chose these two graphs to illustrate two different types of social networks. The Facebook graph represents a network with many clusters, while the Wikipedia graph exemplifies networks centered around a few key nodes. Additionally, to prove that our methods scales to larger graph, we also used the Google+ dataset (Leskovec & Mcauley, 2012) that is constructed from Google circles. It contains 107,614 nodes and 13,673,453 edges.
To conduct experiments on graphs of different sizes from the original, we generated extracted graphs by randomly selecting nodes to match the desired size and then examining the subgraph induced by these nodes. We have verified that selecting random subgraphs does not alter the graph topology of either graphs with respect to the subgraph counting problem.
The focus of our experiments is on assessing accuracy, measured by relative error, as we adjust various parameters. These parameters include download cost, upload cost, graph size, and privacy budget. The results presented are averages calculated from 10 separate simulation runs for each set of parameters. All experiments are comparisons between the following three methods: (1) The leading-edge method detailed in Imola et al. (2022a) (2) Our GroupRR mechanism enhanced by the central server sampling and the clipping method (Subsection 6.4) (3) Our GroupRR mechanism enhanced by the central server sampling and the smooth sensitivity (Subsection 6.3)
7.1.1 Error Analysis for Various Download Costs